OpenStudy (anonymous):

find the ratio of a to b 10 = 2a -- ---- 21 a+b

OpenStudy (aroub):

10(a+b)=2a(21) 10a+10b=42a 10b=32a 32a=10b 32a/b =10 a/b=10/32 a:b=10:32